# Study of multi-particle states with tensor renormalization group method

### v1

@@ 2 Spectroscopy scheme with transfer matrix
where $T$ is the temperature, $s(\bm{r})$ is the spin at $\bm{r}\in\Gamma$, $\hat{\mu}$ is the unit vector for $\mu=0,1$ direction, and $\Gamma$ is a two-dimensional square lattice defined by
$\Gamma=\{\bm{r}=(t,x)|t=0,1,2,\ldots,L_{\rm t}-1\text{~and~}x=0,1,2,\ldots,L_{\rm s}-1\}.$ (2)
Here $L_{\rm t}$ and $L_{\rm s}$ are the system size in time and space direction, respectively. For system under periodic boundary condition (PBC), the partition function can be written as the trace of the transfer matrix
@@
where $\cal{T}$ is explicitly given by
@@
The term inside the first and second parentheses represent the interaction of spin fields in the time and space direction, respectively, while $S,S^{\prime}$ denote the spins configurations at time $t+1$ and $t$
$S^{\prime}$ $=$ $\{s(t+1,x)|x=0,1,2,\ldots,L_{\rm s}-1\},$ (5)
$S$~ $=$ $\{s(t~~~~~~,x)|x=0,1,2,\ldots,L_{\rm s}-1\}.$ (6)
@@
where $\lambda_{a}=e^{-E_{a}}$ is the eigenvalue corresponding to energy $E_{a}$, and $U$ is the unitary matrix of the eigenvectors. Accordingly, the energy gap of the system, $\omega_{a}=E_{a}-E_{\Omega}$, can be obtained from
$\omega_{a}=\log\left(\frac{\lambda_{0}}{\lambda_{a}}\right)\hskip 28.45274pt\mbox{ for }a=1,2,3,\ldots.$ (8)
Here, $E_{\Omega}$ is the ground state energy corresponding to $a=0$.
@@ 3 Tensor network representation and coarse-graining procedure
where $W$ is the unitary matrix and $\lambda_{a}$ is the same eigenvalues as in Eq. (7). In Fig. 1(b), we can see that $\mathcal{A}$ can be obtained by contracting a four-legs tensor $A_{k_{x}l_{x}j_{x}l_{x-1}}$
@@
This four-legs tensor $A_{k_{x}l_{x}j_{x}l_{x-1}}$ is obtained from
@@ 4.1 Energy spectrum and quantum number classification
We compute $\omega_{a}^{[\rm hotrg]}$ in Eq. (18) by coarse graining the pure tensor network with size $L_{\rm t}\times L_{\rm s}=8\times 64$ using HOTRG with a cut-off bond dimension $\chi=80$. We then identify the corresponding quantum number by evaluating $\langle\Omega|\hat{s}_{0}|a\rangle$. We coarse grain the impurity tensor network for $\hat{s}_{0}$ which has the same size as the pure network, and insert the resulting quantities into Eq. (23) to obtain $\langle\Omega|\hat{s}_{0}|a\rangle^{[\rm hotrg]}$. Based on the selection rule given in Eq. (10), if $|\langle\Omega|\hat{s}_{0}|a\rangle^{[\rm hotrg]}|\neq 0$, the quantum number of eigenstate $|a\rangle$ is q=1q=-1; otherwise, the quantum number is $q=+1$. The numerical result of energy with the corresponding quantum number is shown in upper panel of Fig. 3(b). The middle panel shows the results of matrix elements for a=141a=1-41. We observe non-zero values for the eigenstates number a=1,2,3,4,5,7,8,14,15,20,25,26,31,32,33,34,36,37,42a=1,2,3,4,5,7,8,14,15,20,25,26,31,32,33,34,36,37,42, indicating that these states belong to the q=1q=-1 sector, while the remaining states correspond to the $q=+1$ sector. The exact quantum number, computed following Ref. [7], are shown in the bottom panel which we find the good agreement with the numerical results. Using $L_{\rm t}=8$, we are able to compute the energy and identify quantum number up to $a=42$, which is significantly higher than in Ref. [4, 5].
@@ 4.4 Two-particle scattering phase shift
$\delta(k)+kL_{\rm s}/2=0\mod\pi n$ (27)
where $k$ is the relative momentum of the two-particle state, obtained from lattice dispersion relation $k=\cos^{-1}(\cosh m+1-\cosh\omega/2)$, with $\omega_{a}^{[\rm hotrg]}$ is the input energy.
Next we describe how to compute the phase shift by fitting the two-particle state wave function $\psi(x)^{[\rm hotrg]}$ with the free wave function, which is defined as
$\psi(x)^{[\rm free]}=A\cos(k(x+L_{\rm s}/2))\hskip 28.45274pt\text{for~}x>R$ (28)
where $R$ denotes the effective interaction range, defined as the region beyond which the effective potential vanishes. The effective potential is defined as
@@
where $(\partial^{2})^{\rm lat}\psi(x)=\psi(x+1)-2\psi(x)+\psi(x-1)$, and $\psi(x)^{[\rm hotrg]}$ is taken as the input wave function. Note that $k_{\rm lat}^{2}=2(1-\cos{k})$, where $k$ is from the lattice dispersion relation. From Fig. 6(a), we observe that for $T=2.44$, $V_{\rm eff}(x)\to 0$ when $R\approx 40$. Since the fitting method is only applicable when the condition $L_{\rm s}/2>R$ is satisfied, we perform the wave function fitting only for $L_{\rm s}=96,112$. The fitting parameters are $A,k$ and the phase shift is extracted as $\delta(k)=kL_{\rm s}/2\mod{n\pi}$ for $n\in\mathbb{Z}$. In Fig. 6(b), we observe good agreement between the phase shift obtained from the wave function fitting and the finite-volume energy method, as well as consistency with the exact prediction $\delta(k)^{[\rm exact]}=-\pi/2$ [8].
@@ 5 Summary
We have investigated of the multi-particle states of the (1+1)d Ising model using a spectroscopy scheme based on tensor renormalization group method. Our results showed that tensor network with $L_{\rm t}=8$ yields energy spectrum with relatively smaller errors. As a consequence, energies of higher excited states can be reliably computed and the corresponding quantum numbers can be correctly identified. We also correctly identified the eigenstates with zero total momentum in the q=1q=-1 and $q=+1$ sector. From the system size dependence of the energy levels in $q=1,P=0$ sector, we observed one- and three-particle states. Similarly, the two-particle state energies are identified from the energy spectrum in $q=+1,P=0$ sector. Next, we computed the two-particle state wave function using the impurity tensor network method, and successfully extract the two-particle state wave function from the ground to the second excited state. Finally, we extracted the two-particle scattering phase shift using Lüscher’s formula and a fitting procedure applied to the wave function outside the interaction range $R\approx 40$ for $T=2.44$. The phase shift obtained from both methods are consistent with each other and agree with the exact result [8]. As future work, we plan to analyze the three-particle states in detail and apply this scheme to other quantum field theories.
@@ Acknowledgments
F.I.A. is supported by JST SPRING, Grant No. JPMJSP2135. S.T. is supported in part by JSPS KAKENHI Grants No. 21K03531, No. 22H005251, and No. 25K07280. T.Y. is supported in part by JSPS KAKENHI Grant No. 23H01195 and No. 23K25891 , and MEXT as “Program for Promoting Researchers on the Supercomputer Fugaku” (Grant No. JPMXP1020230409).

### v2

@@ 2 Spectroscopy scheme with transfer matrix
where $T$ is the temperature, $s(\bm{r})$ is the spin at $\bm{r}\in\Gamma$, $\hat{\mu}$ is the unit vector for the time ($\mu=0$) and space ($\mu=1$) directions, and $\Gamma$ is a two-dimensional square lattice defined by $\Gamma=\{\bm{r}=(t,x)|t=0,1,2,\ldots,L_{\rm t}-1\text{~and~}x=0,1,2,\ldots,L_{\rm s}-1\}$. Here, $L_{\rm t}$ and $L_{\rm s}$ are the system size in time and space direction, respectively. For system under periodic boundary condition (PBC), the partition function can be written as
@@
where $\cal{T}$ is the transfer matrix, which is explicitly given by
@@
The term inside the first and second parentheses represent the interaction of spin fields in the time and space direction, respectively, while $S^{\prime}=\{s(t+1,x)|x=0,1,2,\ldots,L_{\rm s}-1\}$ and $S=\{s(t,x)|x=0,1,2,\ldots,L_{\rm s}-1\}$ denote the spins configurations at time $t+1$ and $t$.
@@
where $\lambda_{a}=e^{-E_{a}}$ is the eigenvalue corresponding to energy $E_{a}$, and $U$ is the unitary matrix of the eigenvectors. Accordingly, the energy gap of the system, $\omega_{a}=E_{a}-E_{\Omega}$, can be obtained from $\omega_{a}=\log\left(\frac{\lambda_{0}}{\lambda_{a}}\right)$ for $a=1,2,3,\ldots$. Here, $E_{\Omega}$ is the ground state energy corresponding to $a=0$.
@@ 3 Tensor network representation and coarse-graining procedure
where $W$ is the unitary matrix and $\lambda_{a}$ is the same eigenvalues as in Eq. (4). In Fig. 1(b), we can see that $\mathcal{A}$ can be obtained by contracting a four-legs tensor $A_{k_{x}l_{x}j_{x}l_{x-1}}$ [7]
@@
This four-legs tensor $A_{k_{x}l_{x}j_{x}l_{x-1}}$ is obtained from [7]
@@ 4.1 Energy spectrum and quantum number classification
We compute $\omega_{a}^{[\rm hotrg]}$ in Eq. (14) by coarse graining the pure tensor network with size $L_{\rm t}\times L_{\rm s}=8\times 64$ using HOTRG with a cut-off bond dimension $\chi=80$. We then identify the corresponding quantum number by evaluating $\langle\Omega|\hat{s}_{0}|a\rangle$. We coarse grain the impurity tensor network for $\hat{s}_{0}$ which has the same size as the pure network, and insert the resulting quantities into Eq. (19) to obtain $\langle\Omega|\hat{s}_{0}|a\rangle^{[\rm hotrg]}$. Based on the selection rule given in Eq. (6), if $|\langle\Omega|\hat{s}_{0}|a\rangle^{[\rm hotrg]}|\neq 0$, the quantum number of eigenstate $|a\rangle$ is q=1q=-1; otherwise, the quantum number is $q=+1$. The numerical result of energy with the corresponding quantum number is shown in upper panel of Fig. 3(b). The middle panel shows the results of matrix elements for a=142a=1-42. We observe non-zero values for the eigenstates number a=1,2,3,4,5,7,8,14,15,20,25,26,31,32,33,34,36,37,42a=1,2,3,4,5,7,8,14,15,20,25,26,31,32,33,34,36,37,42, indicating that these states belong to the q=1q=-1 sector, while the remaining states correspond to the $q=+1$ sector. The exact quantum number, computed following Ref. [8], are shown in the bottom panel which we find the good agreement with the numerical results. Using $L_{\rm t}=8$, we are able to compute the energy and identify quantum number up to $a=42$, which is significantly higher than in Ref. [4, 5].
@@ 4.4 Two-particle scattering phase shift
$\delta(k)+kL_{\rm s}/2=0\mod\pi n\hskip 28.45274pt\text{for~~}n\in\mathbb{Z}$ (23)
where $k$ is the relative momentum of the two-particle state, obtained from lattice dispersion relation $k=\cos^{-1}(\cosh m+1-\cosh\omega/2)$, with $\omega_{a}^{[\rm hotrg]}$ as the input energy.
Next we describe how to compute the phase shift by fitting the two-particle state wave function $\psi(x)^{[\rm hotrg]}$ with the free wave function, which is defined as $\psi(x)^{[\rm free]}=A\cos(k(x+L_{\rm s}/2))$ for $x>R$, where $R$ denotes the effective interaction range, defined as the region beyond which the effective potential vanishes. The effective potential is defined as
@@
where $(\partial^{2})^{\rm lat}\psi(x)=\psi(x+1)-2\psi(x)+\psi(x-1)$, and $\psi(x)^{[\rm hotrg]}$ is taken as the input wave function. Note that $k_{\rm lat}^{2}=2(1-\cos{k})$, where $k$ is from the lattice dispersion relation. From Fig. 6(a), we observe that for $T=2.44$, $V_{\rm eff}(x)\to 0$ when $R\approx 40$. Since the fitting method is only applicable when the condition $L_{\rm s}/2>R$ is satisfied, we perform the wave function fitting only for $L_{\rm s}=96,112$. The fitting parameters are $A,k$ and the phase shift is extracted as $\delta(k)=kL_{\rm s}/2\mod{n\pi}$. In Fig. 6(b), we observe good agreement between the phase shift obtained from the wave function fitting and the finite-volume energy method, as well as with the exact prediction $\delta(k)^{[\rm exact]}=-\pi/2$ [9].
@@ 5 Summary
We have investigated of the multi-particle states of the (1+1)d Ising model using a spectroscopy scheme based on tensor renormalization group method. Our results showed that tensor network with $L_{\rm t}=8$ yields energy spectrum with relatively smaller errors, enabling reliable determination of higher excited state energies and their corresponding quantum numbers. In particular, we correctly identified eigenstates with zero total momentum in $q=\pm 1$ sectors. From the system size dependence of the energy levels, we identified one- and three-particle states in the q=1,P=0q=-1,P=0 sector, and two-particle states in the $q=+1,P=0$ sector. Next, we computed the two-particle state wave function using the impurity tensor network method, and successfully extract the two-particle state wave function from the ground to the second excited state. Finally, we extracted the two-particle scattering phase shift using Lüscher’s formula and a fitting procedure applied to the wave function outside the interaction range $R\approx 40$ for $T=2.44$. The phase shift obtained from both methods are consistent with each other and agree with the exact result [9]. As future work, we plan to analyze the three-particle states in detail and apply this scheme to other quantum field theories.
@@ Acknowledgments
F.I.A. is supported by JST SPRING, Grant No. JPMJSP2135. S.T. is supported in part by JSPS KAKENHI Grants No. 21K03531, No. 22H05251, and No. 25K07280. T.Y. is supported in part by JSPS KAKENHI Grant No. 23H01195 and No. 23K25891 , and MEXT as “Program for Promoting Researchers on the Supercomputer Fugaku” (Grant No. JPMXP1020230409).